# Analysis of Josephson Junction Barrier Variation - a Combined STEM, Breakdown and Monte-Carlo Approach

### v1

@@ Abstract
Josephson junctions manufactured to tight tolerances are necessary components for superconducting quantum computing. Developing precise manufacturing techniques for Josephson junctions requires an understanding of their make-up and robust feedback metrics against which to optimise. Here we consider complementary techniques assessing what conclusions they allow us to draw about the barriers in junctions. Monte-Carlo simulations of barriers show that standard deviations of 15-20% of the total barrier thickness are compatible with our experimental data. Electrical breakdown allows us to probe the weakest points in barriers. Narrowing the distribution of this breakdown provides a promising feedback mechanism for barrier optimisation. Grouping junctions by breakdown voltage allows us to identify sub-ensembles of junctions with different median resistance, highlighting the importance of weak spots in junction parameters such as resistance. Transmission electron microscopy can be used to find average barrier thickness, although we highlight challenges forming robust conclusions on the distribution of thicknesses in a barrier from these experiments.
@@ I Introduction
Josephson junctions (JJs) are ubiquitous to superconducting quantum computers [1] and other superconducting circuits such as parametric amplifiers [2] and therefore have significant technological importance. The most widely used type of JJ in superconducting technologies is based on an insulating layer of AlOx which separates two superconducting electrodes. The AlOx layer is formed by oxidising the surface of a layer of metallic aluminium before covering the AlOx layer with a second superconducting electrode, also aluminium. This creates a barrier which is (i) amorphous (ii) non-stoichiometric, and (iii) with thickness that varies across the junction [3, 4, 5, 6]. Developing manufacturing techniques for JJs to be made with fine tolerances requires a detailed understanding of the junction, as well as robust feedback metrics against which manufacturing techniques can be optimised.
The figure of merit of an ideal JJ is its critical current, $I_{\rm C}$, a property that can be hard to measure at numbers that are statistically significant. A proxy for this value is the room temperature resistance [7, 8] and resistance tolerances are typically used to characterise manufacturing processes. These processes are in turn optimised considering feedback from resistance measurements and techniques such as scanning electron microscopy [9, 10]. Consequently many problems with repeatable JJ manufacture have been solved and control of the area of JJs is now much improved [7, 11]. As a result the current state of the art in as-fabricated JJ resistance spread is $\sim$2-3[12, 13, 14]. Once lateral dimensions of junctions are perfectly controlled the remaining variation between junctions must be attributed to differences between their barriers although it is not clear what level of resistance variation this causes in typical manufacturing processes.
@@
Variation in dielectric thin films, like the barrier, is a well studied problem with relevance to devices including field effect transistors (FETs) [19], metal-insulator-semiconductor (MIS) or metal-insulator-metal (MIM) structures [20, 21] and magnetic tunnel junctions (MTJs) [22]. The barriers in JJs are typically $\sim 1-2$ nm [23] so even a single O2- ion represents a substantial fraction of the average barrier thickness with ionic radius of $\sim$0.14 nm [24]. Generating a full 3D map of barriers at this resolution isn’t yet possible although various techniques have been used to improve the understanding of barriers. (Scanning) Transmission Electron Microscopy ((S)TEM) and associated techniques give chemical and structural information at sub-nanometre resolution [4, 25]. Atom Probe Tomography (APT) results in 3D reconstructions of the barrier although to date, results on AlOx barriers do not show the requisite resolution [26, 25]. Scanning Tunneling Microscopy (STM) maps the as-grown oxide prior to being capped by a final metal layer [22] and oxygen migration occurs after this final capping [27]. Simulations of the growth process by molecular dynamics give insight into atomic structures in typical barriers [27]. Studies of breakdown voltages of barriers are used in device communities, such as FETs, to probe device-to-device variation in barrier weak points [28] although have not seen substantial use in JJs.
@@ II Electrical measurements
Fig. 1 shows results from fitting a set of 598 IV curves. These are collected from junctions manufactured with the same target dimensions. An atomic force microscope (AFM) measurement of one of these junctions is shown in Fig. 1 (a). For each junction the voltage is swept from 0 V to a value up to 1.6 V as shown in Fig. 1 (b). At low voltages the IVs of the JJs are linear and we fit this dependence at voltages below 0.02 V to extract the junction resistance. At voltages below 1.6 V, each junction fails at its breakdown voltage and becomes an ohmic channel with much lower resistance (typically hundreds of Ohms).
@@
where $A$ is the junction area, $t$ the junction thickness, $\phi$ the barrier height and $K=4\pi t\sqrt{2m_{e}e}/h$. Fitting this model gives a single-valued thickness which, as discussed in the introduction and shown in the literature, for example by STM [22], is a simplification of the true barrier in these junctions. The simple analytical form makes it computationally inexpensive to compute, which is important when used in Monte-Carlo simulations shown later.
[…]
@@
Figure 2: Results from Monte-Carlo simulations of barriers with normally (a - e) and log-normally (f - j) distributed thicknesses. Heatmaps of (a, f) resistance, (b, g) resistance spread (c, h) barrier height ($\phi$), (d, i) fitted thickness and (e, j) fitted area. Resistance and its spread are found by fitting a straight line to the low voltage region and other parameters are from found from fitting Eq. 1 to the full voltage range. Any regions within the experimental variation (or 10% instead for resistance) are coloured grey as shown in the colour bars. Overlaid are ‘bulls-eye’ concentric circle markers, which indicate a set of parameters which satisfies all the constraints set by the experimental measurements.
@@
We characterise each simulated IV curve by refitting it to Eq 1 which allows us to compare simulated IVs to experimental IVs. Fitting the linear regime of the IV gives the resistance of the junction. Repeating this protocol for multiple junctions (20 in this work) gives the junction-to-junction variation for a given thickness distribution. The pixel size is an important parameter in these simulations with smaller pixels giving less junction-to-junction variation. Pixels less than $\sim$0.2 nm become unphysical as they drop below ionic radii of the barrier giving a lower bound on pixel size. We perform these simulations with a pixel size of 1 nm and a barrier height set to the median experimental value shown in Fig. 1 (b) $\phi=1.22$ eV. The pixel size is chosen to correspond approximately to the thickness of barriers assuming that the barrier thickness doesn’t change on a length-scale much less than the thickness and also as a compromise for computational speed. The value of barrier height is important for inferring accurate thicknesses from any subsequent modeling. Literature values vary with changes in barrier height being ascribed to the termination of the oxide barrier (Al/O) and the crystal structure of the underlying metal [30, 31]. Experimental studies also using the Simmons model identify barrier heights ranging from $\sim$0.9 eV to $\sim$1 eV [32]. Density functional theory (DFT) simulations of amorphous Al/AlOx/Al barriers show dependence on roughness of the underlying metal and find effective barrier heights ranging from 0.2 eV to 1 eV [33].
@@
We show the results of these Monte-Carlo simulations in Fig. 2. The top and bottom rows show results for normally distributed and log-normally distributed barrier thicknesses respectively. In each heatmap we colour any cells compatible with our experimental results in grey. For both distributions there is a region where the experimental results shown in Fig. 1 are matched well by simulations indicated by ‘bulls-eye’ symbols in Fig. 2. Using a multi-valued thickness within a junction we are able to match experimental IVs whilst maintaining the known lateral dimensions of the junctions.
@@
Figure 3: Comparison of (a) experimentally measured breakdown voltages with simulated breakdown voltages for (b) log-normally and (c) normally distributed barriers created with distributions indicated by stars in Fig. 2. (a) Shows a double-Gaussian fit indicating a bimodal breakdown distribution. Simulated distributions are found by taking the thinnest points in randomly generated barriers with a mesh size of 0.2 nm and rescaling this by a dielectric strength ($E_{\rm ds}$) to keep the mean breakdown voltage equal to the experimental values. (d) The resistance of junctions distinguishing junctions with breakdown above/below the midpoint of the bimodal distribution 1.3 V, shown by the black dashed line in (a). (e) A histogram of the full thickness distribution of example junctions generated with the parameters from (b, c). (f) Cumulative conductance of junctions from 10 different junctions generated with parameters from (e) as a fraction of the total junction area.
@@
These simulations also show that inferring the barrier height by extracting the value from fits to the Simmons model is not straightforward. In Fig. 2 we use a single value of barrier height $\phi=1.22$ eV. However, the value of $\phi$ extracted from refitting these simulations shown in Fig. 2 (c,h) deviates from this value, depending on the underlying thickness distribution. For instance, moving along the x axis for fixed value of $\langle t\rangle$ in both Fig. 2 (c,h) the barriers with greater thickness variation appear to have a larger barrier heights. This may explain observations such as those seen in Ref. 32 where they conclude that oxidation time has changed the barrier height. Instead of changing barrier height, it may be that the oxidation time is changing thickness distributions (either average or variation), which appear as potential barrier heights when interpreted via fitting to the Simmons model. Given the variation in literature values and imperfect inference from electrical measurements, direct measurements of the barrier height would be advantageous, for instance by x-ray photoelectron spectroscopy techniques such as in Ref. [34].
@@
To understand the effect of changing barrier height, we perform these simulations for $\phi$ = 1 eV and 0.8 eV and show these in the supplemental information and match the results of fitting the Simmons model to data (as indicated by the ‘bulls-eye’ in Fig. 2). With lower barrier heights, we infer a thicker average barrier and also a larger thickness standard deviation. When using the barrier height of 0.8 eV we find average thicknesses which matches our best inference from STEM imaging as discussed below. This places a lower bound on barrier heights compatible with our results. The use of a rectangular potential barrier will give a systematic offset implying a thinner barrier relative to what would be inferred using more realistic rounded potential barriers. A detailed understanding of potential barrier shapes or atomistic computations of tunneling currents would improve this model and likely result in better quantitative agreement between simulated potential barrier thicknesses and the thicknesses of alumina barriers. We could also extend this model to also draw $\phi$ from a distribution of values indicating compositional fluctuations at the expense of extra computational time. An understanding of realistic levels of potential barrier fluctuations would be necessary for this to give deeper insight to our results. The use of a single-valued potential-barrier height here is equivalent to an assumption that fluctuations in thickness contribute more to IVs than fluctuations in barrier height.
@@
We can complement the information shown from the fitted IV curves with measurements of the breakdown voltage, the point indicated by a star in Fig. 1 (a). We present a histogram showing the analysis of junction breakdown voltages of the same group of junctions in Fig. 3 (a). We phenomenologically fit a double-Gaussian to the breakdown voltage showing a bimodal distribution with a midpoint of 1.3 V. Breakdown occurs at the weakest point in the barrier at a voltage given by $V_{\rm BD}=tE_{\rm ds}$ where t is the barrier thickness and $E_{\rm ds}$ its dielectric strength. Assuming a uniform dielectric strength across the barrier, the weakest point is determined by the thinnest point in the junction. Dielectric strength could also fluctuate across the barrier, for instance decreasing at defects in the barrier, where what constitutes a defect in an amorphous oxide is a complicated issue in itself [35]. If all junctions have the same thickness at their thinnest point, then the spread in Fig. 3 (a) would be a direct measure of the spread in dielectric strength, allowing us to place a bound on the standard deviation of this material property in these barriers, $\sigma_{E_{\rm DS}}<4$. We consider the case where dielectric strength is constant across all barriers and within each barrier, i.e. omitting any contributions from local structural changes. In this limit the breakdown voltage is determined by the thinnest point of each barrier and the histogram in Fig. 3 (a) is a rescaled measurement of these thinnest points from 598 junctions.
@@
Extending the Monte-Carlo simulations shown in Fig. 2 we consider whether the same model of varying barrier thickness agrees with our experimental measurements of breakdown voltage. Using thickness distributions identified in Fig. 2 as giving the best match between simulated and experimental IVs we compute a distribution of breakdown voltages. We generate thicknesses from the distributions indicated by ‘bulls-eye’ symbols in Fig. 2 for a $240\times 240$ nm junction with a 0.2 nm pixel size 1110.2 nm approximates the ionic radii in the AlOx barrier. We record the minimum thickness of the barrier in the junction and then repeat this to build a distribution of the thinnest points in the barrier. In order to translate the distribution of the thinnest points to breakdown voltages we rescale the distribution of thin-points by a constant dielectric strength of 3.31 GV/m and 5.26 GV/m for log-normal and normal barrier distributions respectively. Previous measurements of $E_{ds}$ in amorphous alumina vary from 0.4 to 1.0 GV/m , similar in order of magnitude [37]. We compute the dielectric strength by enforcing the mean of the rescaled thicknesses to be equal to the mean of the measured breakdown voltages. Simulated breakdown-distributions for barriers with log-normally and normally distributed thickness are shown in Fig. 3 (b) and (c) respectively. Neither statistical distribution recreates the bimodal distribution seen experimentally. However, the rescaled values from log-normally distributed barriers (Fig. 3 b) show reasonably good agreement with the measured breakdown voltages with a similar standard deviation and min/max ratio. In contrast the rescaled values taken from barriers generated with a normally distributed thickness (Fig. 3 c) do not agree with the experimental measurements. We show that multiple experimental measurements can be largely recreated considering only a skewed thickness distribution in our barriers. The parameters used to simulate breakdown are drawn from simulations of IV curves which have no awareness of breakdown physics and only one value is used to rescale the distribution. We note we do not recreate a bimodal distribution indicating that, while better than a normal distribution, the log-normal thickness distribution is not a perfect description of the barrier.
@@
In Fig. 3 (d) we present histograms of resistance for junctions which have been grouped by the breakdown voltage of the junction, distinguishing junctions as being above or below the mid-point of the bimodal distribution, 1.3 V. We show a difference in average resistance of the high vs. low breakdown junctions of $\sim$300 $\Omega$, a $\sim$4change. This further supports the idea that breakdown occurs at the thin-points in the barrier as anomalously thin points would result in a reduction in resistance and a reduction in breakdown voltage.  [^2] We of course cannot rule out a defect which both reduces the dielectric strength and increases the potential barrier height causing this discrepancy.
@@
In Fig. 3 (e) we show the thickness distribution of two example junctions. We find that the distributions overlap quite well for counts at thinner distributions. Given that the junctions have similar resistances this is good evidence that the majority of the resistance is determined by the thinner portions of the barrier. We see this again in Fig. 3 (f) where we plot a cumulative conductance distribution as a function of junction area and find a large contribution to total conductance coming from a small fraction of the total area, irrespective of whether using a skewed lognormal distribution or a normally distributed thickness. Areas of locally elevated conductance are more prominent in the normally distributed junction, as is to be expected. As the thicker regions contribute little to the total conductance, our model has bad resolution on the thicker-side of the thickness distribution. The key feature of the log-normal probability distribution that our modeling does provide confidence for, is that the tails on the thinner-side of the distribution are suppressed more strongly than in the normal distribution. The observation of a small fraction of the junction contributing a large fraction of conductance agrees qualitatively with the conclusions from Refs. [39, 40]. Both studies arrive at this conclusion from considerations of barrier thickness measurements by (S)TEM, which we consider in the next section.

### v2

# Analysis of Josephson Junction Barrier Variation - a Combined Electron Microscopy, Breakdown and Monte-Carlo Approach

@@ Abstract
Josephson junctions manufactured to tight tolerances are necessary components for superconducting quantum computing. Developing precise manufacturing techniques for Josephson junctions requires an understanding of their make-up and robust feedback metrics against which to optimise. Here we consider complementary techniques assessing what conclusions they allow us to draw about the barriers in junctions. Monte-Carlo simulations of barriers show that standard deviations of 15-20% of the total barrier thickness are compatible with our experimental data. Electrical breakdown allows us to probe the weakest points in barriers. Narrowing the distribution of this breakdown provides a promising feedback mechanism for barrier optimisation. Grouping junctions by breakdown voltage allows us to identify sub-ensembles of junctions with different median resistance. Transmission electron microscopy can be used to find average barrier thickness, although we highlight challenges forming robust conclusions on the distribution of thicknesses in a barrier from these experiments.
@@ I Introduction
Josephson junctions (JJs) are ubiquitous to superconducting quantum computers [1] and other superconducting circuits such as parametric amplifiers [2] and therefore have significant technological importance. The most widely used type of JJ in superconducting technologies is based on an insulating layer of AlOx which separates two superconducting electrodes. The AlOx layer is formed by oxidising the surface of a layer of metallic aluminum before covering the AlOx layer with a second superconducting electrode, also aluminum. This creates a barrier which is (i) amorphous (ii) non-stoichiometric, and (iii) with thickness that varies across the junction [3, 4, 5, 6]. Developing manufacturing techniques for JJs to be made with fine tolerances requires a detailed understanding of the junction, as well as robust feedback metrics against which manufacturing techniques can be optimised.
Typically, when designing superconducting devices reliant on JJs, a specific critical current value, $I_{\rm C}$, is targeted. Ensemble measurements of $I_{\rm C}$ are typically used to characterize Nb/AlOx/Nb JJ properties [7, 8, 9]. This property can be harder to directly measure across statistically significant ensembles in Al/AlOx/Al JJs due to their critical temperature being substantially below liquid helium temperatures. A proxy for this value in junctions with clean insulating barriers is the room temperature resistance [10, 11] and resistance tolerances are typically used to characterise manufacturing processes. These processes are in turn optimised considering feedback from resistance measurements and techniques such as scanning electron microscopy [12, 13]. Consequently many problems with repeatable JJ manufacture have been solved and control of the area of JJs is now much improved [10, 14]. As a result the current state of the art in as-fabricated JJ resistance spread is $\sim$2-3% [15, 16, 17]. Once lateral dimensions of junctions are perfectly controlled the remaining variation between junctions must be attributed to differences between their barriers although it is not clear what level of resistance variation this causes in typical manufacturing processes.
@@
Variation in dielectric thin films, like the barrier, is a well studied problem with relevance to devices including field effect transistors [22], metal-insulator-semiconductor or metal-insulator-metal structures [23, 24] and magnetic tunnel junctions [25]. The barriers in JJs are typically $\sim 1-2$ nm [26] so even a single O2- ion represents a substantial fraction of the average barrier thickness with ionic radius of $\sim$0.14 nm [27]. Generating a full 3D map of barriers at this resolution isn’t yet possible although various techniques have been used to improve the understanding of barriers. (Scanning) Transmission Electron Microscopy ((S)TEM) and associated techniques give chemical and structural information at sub-nanometre resolution [4, 28]. Atom Probe Tomography results in 3D reconstructions of the barrier although to date, results on AlOx barriers do not show the requisite resolution [29, 28]. Scanning Tunneling Microscopy (STM) maps the as-grown oxide prior to being capped by a final metal layer [25] and oxygen migration occurs after this final capping [30]. Simulations of the growth process by molecular dynamics give insight into atomic structures in typical barriers [30]. Studies of breakdown voltages of barriers are used in device communities, such as field effect transistors, to probe device-to-device variation in barrier weak points [31] although have not seen substantial use in JJs.
@@
We study tunnel barriers made by shadow angle evaporation [32] of aluminum forming Al/AlOx/Al tunnel barriers which, when cooled to cryogenic temperatures, form Josephson junctions. Using the same processes as in  [16], junctions are defined in a Dolan bridge geometry  [32] via electron beam lithography. A base layer of aluminum is evaporated onto a sapphire substrate and is then oxidised in a static oxidation process before being capped with a top layer of aluminum and lifted off. All junctions in this work are fabricated on a single sapphire wafer in one fabrication run.
@@ II Electrical measurements
[…]
@@
where $A$ is the junction area (constrained to the value found by AFM in Fig. 1 a), $t$ the junction thickness, $\phi$ the barrier height and $K=4\pi t\sqrt{2m_{e}e}/h$. Fitting this model gives a single-valued thickness which, as discussed in the introduction and shown in the literature, for example by STM [25], is a simplification of the true barrier in these junctions. The simple analytical form makes it computationally inexpensive to compute, which is important when used in Monte-Carlo simulations shown later.
In Fig. 1 (c, d) we show the results of fitting Eq. 1 presenting $t$ and $\phi$ respectively. We find median $\pm$ standard deviation values of the fit parameters given by $t$ = 0.78 nm $\pm 3\%$, $\phi=$1.48 eV $\pm 6\%$. In Fig. 1 (e) we show a histogram of the JJ resistances extracted from the linear fit, centred at 7122 $\Omega$ with a normally distributed resistance and a standard deviation of 3.7 %. We use these experimental results to define targets for Monte-Carlo simulations of IV curves.
@@
Figure 2: Results from Monte-Carlo simulations of barriers with normally (a - e) and log-normally (f - j) distributed thicknesses. Heatmaps of (a, f) resistance, (b, g) resistance spread (c, h) barrier height ($\phi$), (d, i) fitted thickness. Resistance and its spread are found by fitting a straight line to the low voltage region of the simulated IV and other parameters are from found from fitting Eq. 1 to the full voltage range of the simulated IV. Any regions within the experimental variation (or 10% instead for resistance) are shown using a purple to orange heatmap, whereas values outside this range are shown in gray-scale (e,j) For each pixel we sum the number of experimental criteria matched by the simulation. Dark hues indicate regions where the experiment and simulation coincide, indicating that there is a parameter regime for each distribution where this matching is achieved.
@@
We characterise each simulated IV curve by refitting it to Eq 1 which allows us to compare simulated IVs to experimental IVs. In the fitting routine we constrain the area of the junction to be the same value as that found by AFM. Fitting the linear regime of the IV gives the resistance of the junction. Repeating this protocol for multiple junctions (20 in this work) gives the junction-to-junction variation for a given thickness distribution. The pixel size is an important parameter in these simulations with smaller pixels giving less junction-to-junction variation. Pixels less than $\sim$0.2 nm become unphysical as they drop below ionic radii of the barrier [27] giving a lower bound on pixel size. We perform these simulations with a pixel size of 1 nm and a barrier height of $\phi=1.22$ eV. The pixel size is chosen to correspond approximately to the thickness of barriers assuming that the barrier thickness doesn’t change on a length-scale much less than the thickness and also as a compromise for computational speed. The value of barrier height is important for inferring accurate thicknesses from any subsequent modeling. Literature values vary with changes in barrier height being ascribed to the termination of the oxide barrier (Al/O) and the crystal structure of the underlying metal [34, 35]. Experimental studies also using the Simmons model identify barrier heights ranging from $\sim$0.9 eV to $\sim$1 eV [36]. Density functional theory (DFT) simulations of amorphous Al/AlOx/Al barriers show dependence on roughness of the underlying metal and find effective barrier heights ranging from 0.2 eV to 1 eV [37].
@@
We show the results of these Monte-Carlo simulations in Fig. 2. We compare these to experimental results with different criteria for matching to experimental results. Resistance is considered a match if the simulated value is within 10% of experimental values. Resistance spread is considered a match if the spread is $\leq$ 2.4%, as presented in the supplemental information of Ref. [16] for junctions made with this recipe. Simulated values of $t$ and $\phi$ are considered to match if they are within one standard deviation of the median value found by fitting ensembles of experimental IV curves as shown in Fig. 1. The top and bottom rows show results for normally distributed and log-normally distributed barrier thicknesses respectively. In each heatmap we colour any cells compatible with our experimental with a purple - orange heatmap, and show cells incompatible with experimental results in greyscale. For both distributions there is a region where the experimental results shown in Fig. 1 are matched well by simulations indicated by the darkest hues in Fig. 2 (e,j). Using a multi-valued thickness within a junction we are able to match experimental IVs whilst maintaining the known lateral dimensions of the junctions.
@@
Figure 3: Comparison of (a) experimentally measured breakdown voltages with simulated breakdown voltages for (b) log-normally and (c) normally distributed barriers created with distributions in regions where all experimental criteria are matched in Fig. 2. (a) Shows a double-Gaussian fit indicating a bimodal breakdown distribution. Simulated distributions are found by taking the thinnest points in randomly generated barriers with a mesh size of 0.2 nm and rescaling this by a dielectric strength ($E_{\rm ds}$) to keep the mean breakdown voltage equal to the experimental values. (d) The resistance of junctions distinguishing junctions with breakdown above/below the midpoint of the bimodal distribution 1.3 V, shown by the black dashed line in (a). (e) A histogram of the full thickness distribution of example junctions generated with the parameters from (b, c). (f) Cumulative conductance of junctions from 10 different junctions generated with parameters from (e) as a fraction of the total junction area.
@@
These simulations also show that inferring the barrier height by extracting the value from fits to the Simmons model is not straightforward. In Fig. 2 we use a single value of barrier height $\phi=1.22$ eV. However, the value of $\phi$ extracted from refitting these simulations shown in Fig. 2 (c,h) deviates from this value, depending on the underlying thickness distribution. For instance, moving along the x axis for fixed value of $\langle t\rangle$ in both Fig. 2 (c,h) the barriers with greater thickness variation appear to have a larger barrier heights. This may explain observations such as those seen in Ref. 36 where they conclude that oxidation time has changed the barrier height. Instead of changing barrier height, it may be that the oxidation time is changing thickness distributions (either average or variation), which appear as potential barrier heights when interpreted via fitting to the Simmons model. Given the variation in literature values and imperfect inference from electrical measurements, direct measurements of the barrier height would be advantageous, for instance by x-ray photoelectron spectroscopy techniques such as in Ref. [38].
@@
To understand the effect of changing barrier height, we perform these simulations for $\phi$ = 1.5 eV, 1 eV and 0.8 eV and show these in the supplemental information and compare the results of fitting the Simmons model to data. With large barrier heights (similar to the median values found experimentally) we find a parameter regime where simulations match experiment. This parameter regime has very uniform barriers (barriers which vary less than an ionic radius). This is contrary to observations such as those in Ref. [39, 40] where thickness variation can be seen at grain boundaries. For lower barrier heights, we do not find any regions where all parameters match experiments. The parameters which match the most experimentally imposed conditions imply a thicker average barrier and also a larger thickness standard deviation. When using the barrier height of 0.8 eV we find average thicknesses which matches our best inference from STEM imaging discussed below. This places a lower bound on barrier heights compatible with our results. The use of a rectangular potential barrier will give a systematic offset implying a thinner barrier relative to what would be inferred using more realistic rounded potential barriers. A detailed understanding of potential barrier shapes or atomistic computations of tunneling currents would improve this model and likely result in better quantitative agreement between simulated potential barrier thicknesses and the thicknesses of AlOx barriers. We could also extend this model to also draw $\phi$ from a distribution of values indicating compositional fluctuations at the expense of extra computational time. An understanding of realistic levels of potential barrier fluctuations would be necessary for this to give deeper insight to our results. The use of a single-valued potential-barrier height here is equivalent to an assumption that fluctuations in thickness contribute more to IVs than fluctuations in barrier height.
@@
We can complement the information shown from the fitted IV curves with measurements of the breakdown voltage, the point indicated by a star in Fig. 1 (b). We present a histogram showing the analysis of junction breakdown voltages of the same group of junctions in Fig. 3 (a). We phenomenologically fit a double-Gaussian to the breakdown voltage showing a bimodal distribution with a midpoint of 1.3 V. Breakdown occurs at the weakest point in the barrier at a voltage given by $V_{\rm BD}=tE_{\rm ds}$ where t is the barrier thickness and $E_{\rm ds}$ its dielectric strength. Assuming a uniform dielectric strength across the barrier, the weakest point is determined by the thinnest point in the junction. Dielectric strength could also fluctuate across the barrier, for instance decreasing at defects in the barrier, where what constitutes a defect in an amorphous oxide is a complicated issue in itself [41]. If all junctions have the same thickness at their thinnest point, then the spread in Fig. 3 (a) would be a direct measure of the spread in dielectric strength, allowing us to place a bound on the standard deviation of this material property in these barriers, $\sigma_{E_{\rm DS}}<4$ %. We consider the case where dielectric strength is constant across all barriers and within each barrier, i.e. omitting any contributions from local structural changes. In this limit the breakdown voltage is determined by the thinnest point of each barrier and the histogram in Fig. 3 (a) is a rescaled measurement of these thinnest points from 597 junctions.
Extending the Monte-Carlo simulations shown in Fig. 2 we consider whether the same model of varying barrier thickness agrees with our experimental measurements of breakdown voltage. Using thickness distributions identified in Fig. 2 as giving the matches between simulated and experimental IVs, we compute a distribution of breakdown voltages. For Normal (Lognormal) distributions we use a thickness standard deviation and average of 0.21 and 1.025 nm (0.155 and 1 nm) respectively.
@@
We generate thicknesses from the distributions within the region of experimental matching in Fig. 2 for a $240\times 240$ nm junction with a 0.2 nm pixel size 1110.2 nm approximates the ionic radii in the AlOx barrier. We record the minimum thickness of the barrier in the junction and then repeat this to build a distribution of the thinnest points in the barrier. In order to translate the distribution of the thinnest points to breakdown voltages we rescale the distribution of thin-points by a constant dielectric strength of 3.31 GV/m and 5.26 GV/m for log-normal and normal barrier distributions respectively. Previous measurements of $E_{ds}$ in amorphous alumina vary from 0.4 to 1.0 GV/m , similar in order of magnitude [43]. We compute the dielectric strength by enforcing the mean of the rescaled thicknesses to be equal to the mean of the measured breakdown voltages. Simulated breakdown-distributions for barriers with log-normally and normally distributed thickness are shown in Fig. 3 (b) and (c) respectively. Neither statistical distribution recreates the bimodal distribution seen experimentally. However, the rescaled values from log-normally distributed barriers (Fig. 3 b) show reasonably good agreement with the measured breakdown voltages with a similar standard deviation and min/max ratio. In contrast the rescaled values taken from barriers generated with a normally distributed thickness (Fig. 3 c) do not agree with the experimental measurements. We show that multiple experimental measurements can be largely recreated considering only a skewed thickness distribution in our barriers. The parameters used to simulate breakdown are drawn from simulations of IV curves which have no awareness of breakdown physics and only one value is used to rescale the distribution. We note we do not recreate a bimodal distribution indicating that, while better than a normal distribution, the log-normal thickness distribution is not a perfect description of the barrier.
@@
In Fig. 3 (d) we present histograms of resistance for junctions which have been grouped by the breakdown voltage of the junction, distinguishing junctions as being above or below the mid-point of the bimodal distribution, 1.3 V. We show a difference in average resistance of the high vs. low breakdown junctions of $\sim$300 $\Omega$, a $\sim$4% change.
@@
In Fig. 3 (e) we show the thickness distribution of two example simulated junctions. We find that the distributions overlap quite well for counts at thinner distributions. Given that the junctions have similar resistances this is good evidence that the majority of the conductance is determined by the thinner portions of the barrier. We see this again in Fig. 3 (f) where we plot a cumulative conductance distribution as a function of junction area and find a large contribution to total conductance coming from a small fraction of the total area, irrespective of whether using a skewed lognormal distribution or a normally distributed thickness. Areas of locally elevated conductance are more prominent in the normally distributed junction, as is to be expected. As the thicker regions contribute little to the total conductance, our model has bad resolution on the thicker-side of the thickness distribution. The key feature of the log-normal probability distribution that our modeling does provide confidence for, is that the tails on the thinner-side of the distribution are suppressed more strongly than in the normal distribution. The observation of a small fraction of the junction contributing a large fraction of conductance agrees qualitatively with the conclusions from Refs. [44, 40]. Both studies arrive at this conclusion from considerations of barrier thickness measurements by (S)TEM, which we consider in the next section.